OpenStudy (anonymous):

If r(x) = 2 – x2 and w(x) = x – 2, what is the range of (w*r)(x)

OpenStudy (solomonzelman):

is it ` (wº r)(x) `, or is it ` (w • r)(x) ` ?

OpenStudy (anonymous):

the first one

OpenStudy (solomonzelman):

Ok, you know that ` (wº r)(x) ` just means ` w(r(x)) ` in other words, you take the x in the function w(x), and plug in the function r(x) instead of it.

OpenStudy (solomonzelman):

w(x)=x-2 r(x)=2-x² ------------- so, you know that: w(\(\small\color{red}{x}\))=\(\small\color{red}{x}\)-2 and therefore: w(\(\small\color{red}{r(x)}\)) =(\(\small\color{red}{2-x^2}\))-2

OpenStudy (solomonzelman):

can you simplify the w(r(x)) /

OpenStudy (solomonzelman):

?

OpenStudy (anonymous):

do you mean solve (w-x^2)-2?

OpenStudy (anonymous):

2 instead of w

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

if so it would be -x^2

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

so (wº r)(x) = -x² and you need to find the range of this

OpenStudy (anonymous):

ok and how would i do that haha

OpenStudy (solomonzelman):

x² \(\small \ge\) 0 (for any real number value of x)

OpenStudy (solomonzelman):

so the range of x² is (0,+∞)

OpenStudy (solomonzelman):

now, multiply both sides of: x² \(\small \ge\) 0 by -1, and you get

OpenStudy (anonymous):

-x^2 and 0

OpenStudy (solomonzelman):

and what is the sign, \(\small \ge\) or \(\small \le\)

OpenStudy (anonymous):

<

OpenStudy (solomonzelman):

-x² \(\small \le\) 0 like that?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

that is correct

OpenStudy (anonymous):

yay ty

OpenStudy (solomonzelman):

so -x² is at most 0, and can go into any very big negative values

OpenStudy (anonymous):

mhm

OpenStudy (solomonzelman):

so the range of -x² is?

OpenStudy (solomonzelman):

well, it is at most 0, so we know that: ,0] and it can go into nagetive infinty so we know that: (-∞, 0]

OpenStudy (anonymous):

thats it

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

i just didnt know how to make a infinity sign haha

OpenStudy (solomonzelman):

oh, 1. you click ALT, and hold it 2. click 2 3 6 (respectively) on the number pad on the right side of your keyboard 3. you release the ALT, and get ∞

OpenStudy (solomonzelman):

try it yourself

OpenStudy (anonymous):

thank you!

OpenStudy (solomonzelman):

Yes, and if you use a the code 0 2 1 5 , you get ×

OpenStudy (solomonzelman):

and code of 2 4 6 , you get ÷

OpenStudy (solomonzelman):

there are many symbols like that. 755 , ≤ 754, ≥ 251, √

OpenStudy (solomonzelman):

if you have any questions, I will be happy to answer.